OpenStudy (anonymous):

Use the appropriate sum or difference identity to find the exact value of cos (7pi/12)

OpenStudy (anonymous):

cos(7pi/12) = cos(3pi/12+4pi/12) = cos (pi/4 + pi/3) = cos pi/4 * cos pi/3 - sin pi/4 * sin pi/3 go ahead.. put values and get answer :)

OpenStudy (anonymous):

ive got (sqrt2 +sqrt6)/4 is that right